given the equation 2x + 3y =270
solve for y if x = 15
solve for x if y = 15

Answer by rfer(16322)

2x+3y=270
2(15)+3y=270
3y=240
y=80
-------------
2x+3(15)=270
2x+45=270
2x=225
x=112.5